Find the surface area of the portion of the curve x2 =4y from y =1 to y =2 when it is revolved about the y-axis.



1
Expert's answer
2022-02-07T12:27:16-0500

"x^2=4y\quad \Rightarrow \quad x=\sqrt{4y}=2\sqrt{y}"


"g(y)=2\sqrt{y}" from "y=1" to "y=2"

"g^\prime (y)=1/\sqrt{y}"


Surface Area "=\int\limits_1^22\pi \ g(y)\ \sqrt{1+\big(g^\prime (y)\big)^2}\ dy=\int\limits_1^24\pi \sqrt{y} \ \cdot \ \sqrt{1+1/y}\ dy=4\pi \int\limits_1^2\sqrt{y+1}\ dy= \tfrac{8\pi}{3} (y+1)^{3/2}\bigg|_1^2=\tfrac{8\pi}{3}\big(3\sqrt{3}-2\sqrt{2}\big)"


Answer: Surface Area "\approx \ 19.836"